OpenStudy (anonymous):

Let y=2x2. Find the change in y, delta y when x=4 and delta x=03

hartnn (hartnn):

\(\large \Delta y=4x\Delta x\) just plug in values!

OpenStudy (anonymous):

how did you get that?

hartnn (hartnn):

can you differentiate y=2x^2 ?

OpenStudy (anonymous):

yes 4x

hartnn (hartnn):

thats dy/dx = 4x dy = 4x dx

OpenStudy (anonymous):

I see.

OpenStudy (anonymous):

I plug numbers into the equation and got an answer but the online homework says its wrong

hartnn (hartnn):

x=.. ? 3? why u wrote 03 ?

OpenStudy (anonymous):

0.3

OpenStudy (anonymous):

sorry about that

hartnn (hartnn):

delta y = 4*4*0.3 = 16*0.3 =4.8

OpenStudy (anonymous):

I got 4.8 as well but it told me that it was wrong

hartnn (hartnn):

screenshot ?

OpenStudy (anonymous):

sorry dont have a camera right now

OpenStudy (anonymous):

wish i had

hartnn (hartnn):

there must be a PrtSc key in your keyboard ?

OpenStudy (anonymous):

whats that

hartnn (hartnn):

it captures screenshot. that you can paste in Paint and save it

OpenStudy (anonymous):

No i cannot find it! but i found print screen bottom

hartnn (hartnn):

good...press it and use paint...

OpenStudy (anonymous):

ok, i pressed print screen and open paint program

OpenStudy (anonymous):

whats next?

hartnn (hartnn):

paste it or ctrl+v

OpenStudy (anonymous):

this is really great tool! thank you

OpenStudy (anonymous):

I posted it

hartnn (hartnn):

i think i am missing something here.... what we got id dy and not delta y

hartnn (hartnn):

*is

hartnn (hartnn):

for delta y , you use 2(delta x)^2

OpenStudy (anonymous):

,mmmmmm...

hartnn (hartnn):

tried 0.18 for 1st and 4.8 for 2nd ??

OpenStudy (anonymous):

Woop! the second was correct! but not the first one

hartnn (hartnn):

i hope you already tried 4.8 for both ?

OpenStudy (anonymous):

i tried 4.8 for the first but no

hartnn (hartnn):

oh, that is linearization problem!

OpenStudy (anonymous):

yes, it is :)

hartnn (hartnn):

\(y = f(\Delta x )+f'(\Delta x)(x-\Delta x)\)

hartnn (hartnn):

\(y = f(0.3 )+f'(0.3)(4-0.3)\)

hartnn (hartnn):

\(\Delta y= y-f(0.3)\)

hartnn (hartnn):

f' = 4x f'(0.3) = 4*0.3= 1.2 1.2 *(4-0.3) = 1.2 * 3.7=4.44

OpenStudy (anonymous):

is that special equation right? what is it called? the answer is 4.444?

hartnn (hartnn):

that equation is used to make any non-linear equation as approximately linear.

hartnn (hartnn):

only 4.44

OpenStudy (anonymous):

the answer is still wrong

hartnn (hartnn):

1.2*0.3 = 3.6 my last attempt...

hartnn (hartnn):

actually 0.36

OpenStudy (anonymous):

No it didnt work but thanks a lot!